The combustion of octane, C8H18,
proceeds according to the reaction shown.

2C8H18(l)+25O2(g)⟶16CO2(g)+18H2O(l)


Calculate the volume of carbon dioxide produced when 466 moles
of octane combusts at 29.0 °C
and 1.00 bar.
1
Expert's answer
2020-10-14T05:48:14-0400

First of all let's calculate the volume of carbon dioxide produced at STP

stochiometric coefficient ratio of octane and carbon dioxide is 2:16=1:8


466 moles of octane will give 8"\times"466 moles of carbon dioxide

and one mole is equivalent to 22.4 liters at STP

hence total volume at STP = "8\times466\times22.4=83507.2\ liters"


Since STP refers to 0"^o" C and we have to calculate volume at 29o C

"\frac{V_1}{T_1}=\frac{V_2}{T_2}"


"\frac{83507.2}{273}=\frac{V_2}{302}"


"V_2=92377.92liters"